OpenStudy (harmabe91113):

Quadrilateral OPQR is inscribed inside a circle as shown below. What is the measure of angle Q? You must show all work and calculations to receive credit.

OpenStudy (danieldjpon3):

kk

OpenStudy (danieldjpon3):

Can you send the attachment please

OpenStudy (danieldjpon3):

kk brb ima help this kid out real fast

OpenStudy (danieldjpon3):

I'm back

OpenStudy (danieldjpon3):

@harmabe91113

OpenStudy (harmabe91113):

cool

OpenStudy (danieldjpon3):

kk

OpenStudy (danieldjpon3):

im gonna try and solve it

OpenStudy (sshayer):

in a cyclic quadrilateral sum of opposite angles=180

OpenStudy (harmabe91113):

ok thanks

OpenStudy (sshayer):

yw

OpenStudy (danieldjpon3):

92 is teh answer

OpenStudy (danieldjpon3):

And I'll send you my work

OpenStudy (danieldjpon3):

opposite pairs of angles of an inscribed quadrilateral add to 180° O + Q = 180° 2x + 2x + 4 = 180°

OpenStudy (danieldjpon3):

solve for x then substitute into 2x + 4 to get Q

OpenStudy (danieldjpon3):

2x + 2x + 4 = 180

OpenStudy (danieldjpon3):

^ that's the equation you need to solve for x

OpenStudy (danieldjpon3):

After combining you have 4x + 4 = 180 Subtract 4 from both sides

OpenStudy (danieldjpon3):

on the right side left with 174 and on the left we have x

OpenStudy (sshayer):

correct.

OpenStudy (danieldjpon3):

176*

OpenStudy (danieldjpon3):

srry

OpenStudy (danieldjpon3):

so 4x = 176 divide by 4

OpenStudy (harmabe91113):

44

OpenStudy (danieldjpon3):

4x minus 4 is x 180 minus 4 is 176 in other words

OpenStudy (danieldjpon3):

ya 44

OpenStudy (danieldjpon3):

2x + 4, so plug in 44 for x to get Q = 2(44) + 4

OpenStudy (danieldjpon3):

Q = 2x + 4, so plug in 44 for x to get Q = 2(44) + 4

OpenStudy (danieldjpon3):

srry

OpenStudy (danieldjpon3):

this last one is correct

OpenStudy (danieldjpon3):

and u get 92

OpenStudy (danieldjpon3):

and that's the asnwer

OpenStudy (danieldjpon3):

yup

OpenStudy (danieldjpon3):

ok medal pls

OpenStudy (danieldjpon3):

:D

OpenStudy (harmabe91113):

i know :D hold up

OpenStudy (danieldjpon3):

kk (:D

OpenStudy (danieldjpon3):

sure